# Weighted Ensemble Models Are Strong Continual Learners

### v1

@@ 4.1 Experimental Setups
Datasets. Our method was tested on five continuous learning benchmarks. We split each into 10 separate tasks, as done in SLCA [82]. The CIFAR-100 dataset [31] has 100 classes of natural images, each with 500 training images. The ImageNet-R dataset [22] includes images from 200 classes, divided into 24,000 for training and 6,000 for testing. These images, although related to ImageNet-21K, are considered challenging because they are either hard examples from ImageNet or new images in different styles, making them tough for the pre-trained model to adapt. The CUB-200 dataset [68] consists of images from 200 bird classes, with about 60 images per class, half for training and half for testing. The Cars-196 dataset [29] is made up of 196 types of car images, split into 8,144 for training and 8,040 for testing, maintaining a similar class ratio. Lastly, VTAB-50 [81] dataset contains diverse classes from multiple complex realms. We sample 5 datasets from VTAB, each containing 10 classes, to construct the cross-domain CIL setting. The first two datasets are broader in classification, similar to the ones used in [73, 72], while the last three focus on more specific classifications. More details are reported in the Supp.
@@
We compute the average classification accuracy (denoted as Inc-Acc (%)) of all the classes ever seen after learning each incremental task and present the accuracy after learning the last task (denoted as Last-Acc (%)).
@@
Baselines and Competitors: We first compare to the state-of-the-art PTM-based CIL methods L2P [73], DualPrompt [72] and SLCA [82]. We also utilize the same PTM as the initialization for classical CIL methods GDumb [49], LwF [34], DER [4], BIC [77], EWC [28]. Additionally, we report the baselines, sequentially fine-tuning of the model, denoted as Seq FT, and Prototype-classifier is a cosine similarity classifier on the extracted features of the pre-trained model. Joint-Training is an upper bound, where the model has been trained on all the tasks at the same time.
Variant Init. $\bm{\lambda}$ Split CIFAR-100 Split Imagenet-R Split CUB-200 Split Cars-196 Split VTAB-50
Last-Acc Inc-Acc Last-Acc Inc-Acc Last-Acc Inc-Acc Last-Acc Inc-Acc Last-Acc Inc-Acc
Seq FT - - 88.86±$\pm 0.83$ 92.01±$\pm 1.71$ 71.80±$\pm 1.45$ 76.84±$\pm 1.26$ 68.07±$\pm 1.09$ 79.04±$\pm 1.69$ 49.74±$\pm 1.25$ 62.83±$\pm 2.16$ 68.48±$\pm 0.63$ 84.37±$\pm 0.35$
SLCA - - 91.53±$\pm 0.28$ 94.09±$\pm 0.87$ 77.00±$\pm 0.33$ 81.17±$\pm 0.64$ 84.71±$\pm 0.40$ 90.94±$\pm 0.68$ 67.73±$\pm 0.85$ 76.93±$\pm 1.21$ 92.52±$\pm 0.61$ 93.54±$\pm 1.08$
Weight-Ens. $\theta_{0}^{*}$ $1/t$ 61.68±$\pm 0.14$ 70.24±$\pm 0.46$ 44.90±$\pm 0.14$ 52.03±$\pm 0.46$ 82.49±$\pm 0.27$ 87.20±$\pm 0.56$ 36.20±$\pm 0.27$ 45.31±$\pm 0.31$ 89.12±$\pm 0.35$ 89.17±$\pm 0.14$
Weight-Ens. $\theta^{*}_{t-1}$ $1/t$ 91.69±$\pm 0.23$ 94.52±$\pm 0.95$ 75.85±$\pm 0.73$ 81.51±$\pm 0.60$ 84.28±$\pm 0.47$ 90.07±$\pm 0.22$ 71.82±$\pm 0.47$ 78.85±$\pm 0.31$ 91.26±$\pm 0.31$ 91.72±$\pm 0.18$
EMA - - 91.40±$\pm 0.12$ 93.89±$\pm 0.59$ 79.35±$\pm 0.83$ 83.07±$\pm 0.94$ 84.99±$\pm 0.49$ 90.84±$\pm 0.74$ 65.41±$\pm 0.18$ 73.68±$\pm 0.46$ 93.35±$\pm 0.71$ 92.87±$\pm 0.46$
CoMA $\theta^{*}_{t-1}$ $0.5$ 92.00±$\pm 0.13$ 94.12±$\pm 0.63$ 79.32±$\pm 0.05$ 82.32±$\pm 0.17$ 86.95±$\pm 0.29$ 91.75±$\pm 0.39$ 73.35±$\pm 0.59$ 78.55±$\pm 0.42$ 92.23±$\pm 0.52$ 92.30±$\pm 0.43$
CoFiMA (Full) $\theta^{*}_{t-1}$ $0.5F_{t}$ 92.16±$\pm 0.24$ 94.66±$\pm 0.94$ 79.48±$\pm 0.26$ 83.28±$\pm 0.56$ 87.02±$\pm 0.56$ 92.03±$\pm 0.69$ 73.46±$\pm 0.64$ 79.35±$\pm 0.96$ 93.67±$\pm 0.23$ 94.96±$\pm 0.11$
Table 3: Experimental results comparing our methods (CoMA, and CoFiMA) to weight-averaging baselines using ViT-B-16 [11].
@@
Implementation. We adopt a pre-trained ViT-B/16 backbone for all the baselines unless mentioned otherwise. We follow the implementation of SLCA [82] that adapts a small learning rate of 0.0001 for the representation layer and 0.01 for the classification layer with Class-Alignment strategy (CA) that improves the classification layer by modeling the class-wise distributions and aligning the classification layers. We use a batch size of 128 with our hyperparameter $\lambda\!=\!0.5$ in all our experiments.
4.2 Experimental Results
Performance Analysis. This section analyzes the performance of CoFiMA across various continual learning benchmarks including Split CIFAR-100, Split ImageNet-R, Split CUB-200, and Split Cars-196, as detailed in Tabs 1 and 2.
In supervised pre-training (Tab. 1), CoFiMA achieves a Last-Acc of 92.16% and an Inc-Acc of 94.66% on Split CIFAR-100. This performance surpasses SLCA’s performance, demonstrating a gain of 0.63% and 0.57% in Last-Acc and Inc-Acc, respectively. Similarly, on Split ImageNet-R, CoFiMA records a Last-Acc of 79.48%, which is a 2.48% improvement over SLCA’s 77.00%, and an Inc-Acc of 83.28%, outperforming SLCA’s 81.17%. This trend of outperforming SLCA is also observed in Split Cars-196 with a 5.73% improvement.
In self-supervised pre-training scenarios (Tab. 2), CoFiMA maintains its lead. On Split ImageNet-R, CoFiMA’s Last-Acc of 68.15% and Inc-Acc of 73.66% continue to show an advantage over SLCA. CoFiMA also leads in Split CUB-200 and Split Cars-196, showcasing its consistent performance across different datasets.
@@ 4.2 Experimental Results
Notably, CoFiMA’s performance is not far from the Joint-Training baselines. For instance, in Split CIFAR-100 (supervised), CoFiMA’s Last-Acc is only 1.06% lower than the Joint-Training baseline of 93.22%. This gap narrows further in other datasets, indicating CoFiMA’s effectiveness in approaching the upper bounds of CL performance. CoFiMA’s approach, which efficiently balances the retention of old knowledge with the acquisition of new information, contributes to its strong performance in both supervised and self-supervised settings.
Model Averaging. This section evaluates our continual model-averaging approach against two baselines: (i) Weight-Ensemble, which uniformly averages model weights (e.g. $\lambda=1/t$), initializing each model $M_{t}$ from either the pre-trained PTM ($\theta_{0}$) or the previous task’s parameters ($\theta_{t-1}$); (ii) Exponential Moving Average (EMA) [62], a technique for a running average of model parameters. We use debiased EMA such that for each iteration $m\in{1,\ldots,M}$, the model is updated as follows: $\theta_{m}=\beta\theta_{m}+(1-\beta\theta_{m-1})$ with $\beta=0.999$ (see details in Supp.). Tab. 3 presents the results.
@@
CoFiMA shows superior performance across all datasets. The Weight-Ensemble with $\theta_{t-1}$ initialization yields competitive results in Split CIFAR-100 (Last-Acc: 91.69%) but underperforms in datasets like Split Cars-196 (Last-Acc: 71.82%). The Weight-Ensemble starting from $\theta_{0}$ shows lower performance. This indicates the limitations of re-initialization to $\theta_{0}$, as models are trained on different tasks it leads to different optima, as depicted in Fig. 2(a).
@@
However, Weight-Ensemble with $\theta_{t\!-\!1}$ initialization outperforms the variant starting from $\theta_{0}$, emphasizing the importance of initialization. Initializing $\theta_{t}$ from $\theta^{*}_{t\!-\!1}$ improves performance on both task $t$ and $t\!-\!1$, as done in  [75, 24]. This gain in performance indicates that averaging successive models leads to good performance on the current task $t$, while preserving previous knowledge.
@@
Averaging all models trained up to task $t$ is suboptimal, as weights at task $t$ likely differ significantly from those at task $t\!=\!1$. Such averaging shifts the $\theta^{*}_{t}$ values toward a less optimal minima, leading to decreased performance (refer to Fig. 2(b)). In contrast, CoFiMA avoids this by averaging only with the preceding parameters $\theta_{t-1}$ leading to better performance, in line with findings that parameters of over-parametrized pre-trained models undergo minimal change during finetuning [6].
@@
The EMA method, though superior to both Weight-Ensemble variants in most scenarios, falls short of CoFiMA’s performance. This discrepancy may stem from excessive averaging in EMA, where $\theta^{*}_{t}$ is modified multiple times within the same task, potentially leading to suboptimal outcomes. Differently, our method applies weight averaging only after completing each task, enhancing computational efficiency by selectively focusing on pertinent model weights, while retaining previous knowledge.
4.3 Ablation Study
@@
Effect of PTMs. In this section, we evaluate the performance of CoFiMA approach across a variety of backbone architectures, including self-supervised (MAE [21], MoCoV3 [5], and DINOv2 [45]) and supervised (ViT-Tiny [11], and ViT-B/16-SAM [15]) models. This comprehensive analysis aimed to ascertain the adaptability and performance consistency of Co-FiMA in diverse training paradigms. Results are visualized in Fig. 3.
@@
Our results indicate that CoFiMA consistently enhances performance across all tested backbone architectures relative to the baseline SLCA method. For instance, with the ViT-Tiny backbone, Co-FiMA improves the Top-1 Average Accuracy from 80.25% to 82.96%. This trend is similarly observed with ViT-Large, where Co-FiMA achieves an accuracy of 86.81%, surpassing SLCA’s 85.93%.
@@
In the context of self-supervised learning models, CoFiMA demonstrates its efficacy by outperforming the SLCA on the ViT-B/16-DINOv2, ViT-B/16-MAE, and ViT-B/16-MoCoV3 backbones. Importantly, the ViT-B/16-SAM backbone achieves the best performance among the evaluated models. This can be attributed to the effective generalization features ingrained in the SAM backbone, a result of being trained with the SAM optimizer [15]. This optimizer is known for its capability to enhance model generalizability, which is reflected in the superior performance metrics observed in our experiments, also mentioned in Mehta et al. [40] work. We also notice that self-supervised pre-training usually results in larger performance gaps between continual learning baselines and joint training.
When compared to Joint-Training, a standard upper-bound in CL setting, CoFiMA demonstrates competitive performance. Although Joint-Training occasionally leads, as seen with ViT-Large (94.45%) and ViT-B/16-SAM (91.87%), CoFiMA remains close, particularly with ViT-B/16-SAM, where it achieves 90.48%.
@@
These results indicate that CoFiMA is a versatile approach, effective in enhancing the performance of various vision transformer backbones in both supervised and self-supervised learning contexts. However, the performance boost from our approach varies with the choice of backbone (size), and its pertaining paradigm. The consistent performance improvements across different models highlight its potential effectiveness in CL.
Figure 3: Assessing performance of CoFiMA using various PTMs on CIFAR-100 dataset. CoFiMA enhances the results of SLCA. Additionally, it is noted that CoFiMA’s performance is contingent on the characteristics of the chosen PTM backbone.
Balancing the Information of an Old and a New Task.

### v3

@@ 4.1 Experimental Setups
Datasets and Settings. We conduct experiments on four CIL benchmarks: CIFAR-100 [30], ImageNet-R [19], CUB-200 [71], and Cars-196 [28]. The CIFAR-100 dataset [30] has 100 classes of natural images, each with 500 training images. The ImageNet-R dataset [19] includes images from 200 classes, divided into 24,000 for training and 6,000 for testing. These images, although related to ImageNet-21K, are considered challenging for the PTM because they are either hard examples from ImageNet or new images in different styles. The CUB-200 dataset [71] consists of images from 200 bird classes, with about 60 images per class, half for training and half for testing. The Cars-196 dataset [28] is made up of 196 types of car images, split into 8,144 for training and 8,040 for testing, maintaining a similar class ratio. The first two focus on fine-grained classifications, while the last two datasets (i.e., CIFAR-100, and ImageNet-R) are standard benchmarks for CL. Following SLCA [85] we split each benchmark into 10 tasks. We report the results in the class-incremental setting, i.e. the task id is not known during inference.
@@
Metrics. We report the average classification accuracy of all the classes ever seen after learning each incremental task (denoted as Inc-Acc (%)) and the accuracy after learning the last task (denoted as Last-Acc (%)).
@@
Baselines and Competitors. We compare with the state-of-the-art PTM-based CIL methods L2P [78], DualPrompt [77], SLCA [85], and RanPAC [38]. We also utilize the same PTM as the initialization for classical CL methods GDumb [50], LwF [33], DER [4], BIC [82], and EWC [26]. Additionally, we report the following baselines: sequentially fine-tuning of the model (denoted as Seq FT), and Prototype-classifier [23], which is a cosine similarity classifier on the extracted features of the PTM. Joint-Training is an upper bound, where the model has been trained on all the tasks at the same time.
@@
Implementation details. We adopted two kinds of PTMs in our experiments: a ViT-B/16 [27] backbone supervisedly pre-trained on ImageNet-21K [55], the default PTM unless otherwise stated; and ViT-B/16 backbone with self-supervised pre-training using MoCo-V3 [5] on ImageNet-1K. We follow the implementation of SLCA [85] that adapts a small learning rate of 0.0001 for the representation layer and 0.01 for the classification layer, along with the class-alignment strategy. We set the batch size to 128, $\lambda\!=\!0.4$ for supervised and $\lambda\!=\!0.2$ for self-supervised pre-training, in all our experiments.
Table 1: State-of-the-art comparison on CUB-200, Cars-196, CIFAR-100, and ImageNet-R using ViT-B/16 [27] supervisedly pre-trained on ImageNet-21K [55].
Method Memory CUB-200 Cars-196 CIFAR-100 ImageNet-R
Joint-Training - 88.00±$\pm 0.34$ 80.31±$\pm 0.13$ 93.22±$\pm 0.16$ 79.60±$\pm 0.87$
Prototype-classifier [23] 80.66±$\pm 0.00$ 88.95±$\pm 0.00$ 28.58±$\pm 0.00$ 39.83±$\pm 0.00$ 60.29±$\pm 0.00$ 69.18±$\pm 0.00$ 38.45±$\pm 0.00$ 45.59±$\pm 0.00$
GDumb [50] 61.80±$\pm 0.77$ 79.76±$\pm 0.18$ 25.20±$\pm 0.84$ 49.48±$\pm 0.74$ 81.92±$\pm 0.15$ 89.46±$\pm 0.94$ 24.23±$\pm 0.35$ 43.48±$\pm 0.49$
DER++ [4] 77.42±$\pm 0.71$ 87.61±$\pm 0.09$ 60.41±$\pm 1.76$ $\pm 0.57$ 84.50±$\pm 1.67$ 91.49±$\pm 0.61$ 67.75±$\pm 0.93$ 78.13±$\pm 1.14$
BiC [82] 81.91±$\pm 2.59$ 89.29±$\pm 1.57$ 63.10±$\pm 5.71$ 73.75±$\pm 2.37$ 88.45±$\pm 0.57$ 93.37±$\pm 0.32$ 64.89±$\pm 0.80$ 73.66±$\pm 1.61$
L2P [78] $\checkmark$ 62.21±$\pm 1.92$ 73.83±$\pm 1.67$ 38.18±$\pm 2.33$ 51.79±$\pm 4.19$ 82.76±$\pm 1.17$ 88.48±$\pm 0.83$ 66.49±$\pm 0.40$ 72.83±$\pm 0.56$
DualPrompt [77] $\checkmark$ 66.00±$\pm 0.57$ 77.92±$\pm 0.50$ 40.14±$\pm 2.36$ 56.74±$\pm 1.78$ 85.56±$\pm 0.33$ 90.33±$\pm 0.33$ 68.50±$\pm 0.52$ 72.59±$\pm 0.24$
EWC [26] $\checkmark$ 68.32±$\pm 2.64$ 79.95±$\pm 2.28$ 52.50±$\pm 3.18$ 64.01±$\pm 3.25$ 89.30±$\pm 0.23$ 92.31±$\pm 1.66$ 70.27±$\pm 1.99$ 76.27±$\pm 2.13$
LwF [33] $\checkmark$ 69.75±$\pm 1.37$ 80.45±$\pm 2.08$ 49.94±$\pm 3.24$ 63.28±$\pm 1.11$ 87.99±$\pm 0.05$ 92.13±$\pm 1.16$ 67.29±$\pm 1.67$ 74.47±$\pm 1.48$
Seq FT $\checkmark$ 68.07±$\pm 1.09$ 79.04±$\pm 1.69$ 49.74±$\pm 1.25$ 62.83±$\pm 2.16$ 88.86±$\pm 0.83$ 92.01±$\pm 1.71$ 71.80±$\pm 1.45$ 76.84±$\pm 1.26$
RanPAC [38] 85.82±$\pm 0.53$ 91.47±$\pm 0.96$ 53.84±$\pm 0.84$ 66.39±$\pm 1.18$ 90.09±$\pm 0.25$ 93.31±$\pm 0.98$ 72.62±$\pm 0.11$ 78.35±$\pm 0.58$
SLCA [85] $\checkmark$ 84.71±$\pm 0.40$ 90.94±$\pm 0.68$ 67.73±$\pm 0.85$ 76.93±$\pm 1.21$ 91.53±$\pm 0.28$ 94.09±$\pm 0.87$ 77.00±$\pm 0.33$ 81.17±$\pm 0.64$
CoMA (Ours) 85.95±$\pm 0.29$ 90.75±$\pm 0.39$ 73.35±$\pm 0.59$ 78.55±$\pm 0.42$ 92.00±$\pm 0.13$ 94.12±$\pm 0.63$ 77.47±$\pm 0.05$ 81.32±$\pm 0.17$
CoFiMA (Ours) 87.11±$\pm 0.56$ 91.87±$\pm 0.69$ 76.96±$\pm 0.64$ 82.65±$\pm 0.96$ 92.77±$\pm 0.24$ 94.89±$\pm 0.94$ 78.25±$\pm 0.26$ 81.48±$\pm 0.56$
4.2 State-of-the-art Comparison
This section analyzes the performance of CoMA and CoFiMA across various CL benchmarks. We report in Tabs. 1 and 2 the comparison of CoMA and CoFiMA with state-of-the-art CL methods using the ViT-B/16 backbone that was pre-trained supervisedly and unsupervisedly, respectively.
As shown in Tab. 1, our proposed CoMA consistently outperforms the best-performing CIL baselines, SLCA [85] and RanPAC [38], across all benchmarks. This confirms the benefits of model averaging in CIL. Moreover, CoFiMA, the improved variant of CoMA, shows even further performance improvement over CoMA, achieving new state-of-the-art results in CIL. This highlights the need for adaptive model averaging based on the importance of the parameters for a given task. In detail, CoFiMA achieves a Last-Acc of 87.11% and an Inc-Acc of 91.87% on CUB-200. This performance surpasses SLCA’s performance, demonstrating a gain of +2.4% and +0.93% in Last-Acc and Inc-Acc, respectively. Similarly, on Cars-196, CoFiMA outperforms SLCA with +9.23% and +5.72% improvements in Last-Acc and Inc-Acc, respectively. This trend of outperforming SLCA is also observed in Imagenet-R, with +1.25% and +0.31% improvements in Last-Acc and Inc-Acc over SLCA, respectively.
Table 2: State-of-the-art comparison on CUB-200, Cars-196, CIFAR-100, and ImageNet-R using ViT-B/16 with self-supervised pre-training (MoCo-V3 [5]) on ImageNet-1K.
Method Memory CUB-200 Cars-196 CIFAR-100 ImageNet-R
Joint-Training - 79.55±$\pm 0.04$ 74.52±$\pm 0.09$ 89.11±$\pm 0.06$ 72.80±$\pm 0.23$
Prototype-classifier [23] 51.57±$\pm 0.00$ 63.43±$\pm 0.00$ 20.97±$\pm 0.00$ 30.10±$\pm 0.00$ 73.50±$\pm 0.00$ 81.75±$\pm 0.00$ 37.60±$\pm 0.00$ 44.95±$\pm 0.00$
GDumb [50] 45.29±$\pm 0.97$ 66.86±$\pm 0.63$ 20.95±$\pm 0.42$ 45.40±$\pm 0.66$ 69.72±$\pm 0.20$ 80.95±$\pm 1.19$ 28.24±$\pm 0.58$ 43.64±$\pm 1.05$
DER++ [4] 61.47±$\pm 0.32$ 77.15±$\pm 0.61$ 50.64±$\pm 0.70$ 67.64±$\pm 0.45$ 63.64±$\pm 1.30$ 79.55±$\pm 0.87$ 53.11±$\pm 0.44$ 65.10±$\pm 0.91$
BiC [82] 74.39±$\pm 1.12$ 82.13±$\pm 0.33$ 65.57±$\pm 0.93$ 73.95±$\pm 0.29$ 80.57±$\pm 0.86$ 89.39±$\pm 0.33$ 57.36±$\pm 2.68$ 68.07±$\pm 0.22$
EWC [26] $\checkmark$ 61.36±$\pm 1.43$ 72.84±$\pm 2.18$ 53.16±$\pm 1.45$ 63.61±$\pm 1.06$ 81.62±$\pm 0.34$ 87.56±$\pm 0.97$ 64.50±$\pm 0.36$ 70.37±$\pm 0.41$
LwF [33] $\checkmark$ 61.66±$\pm 1.95$ 73.90±$\pm 1.91$ 52.45±$\pm 0.48$ 63.87±$\pm 0.31$ 77.94±$\pm 1.00$ 86.90±$\pm 0.90$ 60.74±$\pm 0.30$ 68.55±$\pm 0.65$
Seq FT $\checkmark$ 61.67±$\pm 1.37$ 73.25±$\pm 1.83$ 52.91±$\pm 1.61$ 63.32±$\pm 1.31$ 81.47±$\pm 0.55$ 87.55±$\pm 0.95$ 64.43±$\pm 0.44$ 70.48±$\pm 0.54$
RanPAC [38] $\checkmark$ 74.43±$\pm 0.43$ 83.63±$\pm 0.01$ 63.21±$\pm 0.02$ $\pm 0.47$ 86.47±$\pm 0.52$ 90.81±$\pm 1.05$ 69.11±$\pm 0.69$ 75.20±$\pm 0.34$
SLCA [85] $\checkmark$ 73.01±$\pm 0.16$ 82.13±$\pm 0.34$ 66.04±$\pm 0.08$ 72.59±$\pm 0.04$ 85.27±$\pm 0.08$ 89.51±$\pm 1.04$ 68.07±$\pm 0.21$ 73.04±$\pm 0.56$
CoMA (Ours) $\checkmark$ 75.12±$\pm 0.27$ 82.76±$\pm 0.16$ 67.48±$\pm 0.19$ 74.90±$\pm 0.87$ 86.59±$\pm 0.51$ 91.02±$\pm 0.47$ 69.33±$\pm 0.22$ 75.64±$\pm 0.13$
CoFiMA (Ours) 77.65±$\pm 0.18$ 83.54±$\pm 0.16$ 69.51±$\pm 0.16$ 76.21±$\pm 0.83$ 87.44±$\pm 0.47$ 91.13±$\pm 0.53$ 70.87±$\pm 0.31$ 76.09±$\pm 0.78$
From Tab. 2 we observe that CoMA and CoFiMA both surpass the state-of-the-art methods by substantial margins while using a PTM pre-trained unsupervisedly on ImageNet-1K. The results confirm that having access to an unsupervised PTM is sufficient to reach satisfactory performance in CIL, although the absolute performance is lower for all methods compared to Tab. 1. In detail, on CUB-200, CoFiMA’s Last-Acc of 77.65% and Inc-Acc of 83.54% continue to show an advantage over SLCA. CoFiMA also leads in Cars-196, CIFAR-100, and ImageNet-R, showcasing its consistent performance across different datasets.
@@ 4.2 State-of-the-art Comparison
Notably, CoFiMA’s performance is not far from the joint-training baselines in both Tabs. 1 and 2. For instance, in Tab. 1 for CIFAR-100, CoFiMA’s Last-Acc is only 0.45% lower than the joint-training baseline of 93.22%. This gap narrows further in other benchmarks, indicating CoFiMA’s effectiveness in approaching the upper bounds of CL performance. CoFiMA, which efficiently balances the retention of old knowledge with the acquisition of new information, contributes to its strong performance using both supervised and unsupervised PTMs.
4.3 Ablation Studies
Analysis of Model Averaging. This section evaluates our continual model-averaging approach against two baselines:
  • Weight-Ensemble, which uniformly averages model weights (e.g. $\lambda\!=\!1/t$), initializing each model $M_{t}$ from either the pre-trained PTM ($\boldsymbol{\theta}_{0}$) or the previous task’s parameters ($\boldsymbol{\theta}_{t-1}$).
  • Exponential Moving Average (EMA) [66], a technique for a running average of model parameters computed at every gradient descent iteration $m$ as follows: $\boldsymbol{\theta}_{m}\!=\!\beta\boldsymbol{\theta}_{m}+(1-\beta)\boldsymbol{\theta}_{m\!-\!1}$ with $\beta=0.999$.
Table 3: Experimental results comparing our methods (CoMA, and CoFiMA) to weight-averaging baselines using ViT-B/16 [27] supervised PTM.
Variant Init. $\boldsymbol{\lambda}$ CUB-200 Cars-196 CIFAR-100 Imagenet-R
Last-Acc Inc-Acc Last-Acc Inc-Acc Last-Acc Inc-Acc Last-Acc Inc-Acc
Prototype-classifier [23] - - 80.66±$\pm 0.00$ 88.95±$\pm 0.00$ 28.58±$\pm 0.00$ 39.83±$\pm 0.00$ 60.29±$\pm 0.00$ 69.18±$\pm 0.00$ 38.45±$\pm 0.00$ 45.59±$\pm 0.00$
Seq FT - - 68.07±$\pm 1.09$ 79.04±$\pm 1.69$ 49.74±$\pm 1.25$ 62.83±$\pm 2.16$ 88.86±$\pm 0.83$ 92.01±$\pm 1.71$ 71.80±$\pm 1.45$ 76.84±$\pm 1.26$
SLCA [85] - - 84.71±$\pm 0.40$ 90.94±$\pm 0.68$ 67.73±$\pm 0.85$ 76.93±$\pm 1.21$ 91.53±$\pm 0.28$ 94.09±$\pm 0.87$ 77.00±$\pm 0.33$ 81.17±$\pm 0.64$
Weight-Ens. $\theta_{0}$ $1/t$ 82.49±$\pm 0.27$ 87.20±$\pm 0.56$ 36.20±$\pm 0.27$ 45.31±$\pm 0.31$ 61.68±$\pm 0.14$ 70.24±$\pm 0.46$ 44.90±$\pm 0.14$ 52.03±$\pm 0.46$
Weight-Ens. $\theta^{*}_{t-1}$ $1/t$ 84.28±$\pm 0.47$ 90.07±$\pm 0.22$ 71.82±$\pm 0.47$ 78.85±$\pm 0.31$ 91.69±$\pm 0.23$ 94.52±$\pm 0.95$ 75.85±$\pm 0.73$ 81.51±$\pm 0.60$
EMA - - 84.99±$\pm 0.49$ 90.84±$\pm 0.74$ 65.41±$\pm 0.18$ 73.68±$\pm 0.46$ 91.40±$\pm 0.12$ 93.89±$\pm 0.59$ 77.35±$\pm 0.83$ 83.07±$\pm 0.94$
[…]
CoFiMA (Ours) $\theta^{*}_{t-1}$ $\lambda F_{t}$ 87.11±$\pm 0.56$ 91.87±$\pm 0.69$ 76.96±$\pm 0.64$ 82.65±$\pm 0.96$ 92.77±$\pm 0.24$ 94.89±$\pm 0.94$ 78.25±$\pm 0.26$ 81.48±$\pm 0.56$
@@ 4.3 Ablation Studies
Tab. 3 presents the results. CoFiMA shows superior performance across all datasets. The Weight-Ensemble with $\theta_{t-1}$ initialization yields competitive results in CIFAR-100 (Last-Acc: 91.69%) but underperforms in datasets like Cars-196 (Last-Acc: 71.82%). The Weight-Ensemble starting from $\theta_{0}$ shows lower performance. This indicates the limitations of reinitialization to $\theta_{0}$; as models are trained on different tasks, it leads to different optima, as depicted in Fig. 2(a).
@@
However, Weight-Ensemble with $\boldsymbol{\theta}_{t\!-\!1}$ initialization outperforms the variant starting from $\boldsymbol{\theta}_{0}$, emphasizing the importance of initialization. Initializing $\boldsymbol{\theta}_{t}$ from $\boldsymbol{\theta}^{*}_{t\!-\!1}$ improves performance on both task $t$ and $t\!-\!1$, as done in [80, 21]. This gain in performance indicates that averaging successive models leads to good performance on the current task $t$ while preserving previous knowledge.
@@
Averaging all models trained up to task $t$ is suboptimal, as weights at task $t$ likely differ significantly from those at task $t\!=\!1$. Such averaging shifts the $\boldsymbol{\theta}^{*}_{t}$ values toward suboptimal minima, leading to decreased performance (refer to Fig. 2(b)). In contrast, CoFiMA avoids this by averaging only with the preceding parameters $\boldsymbol{\theta}^{*}_{t-1}$, leading to better performance.
@@
The EMA method, though superior to both Weight-Ensemble variants in most scenarios, falls short of CoFiMA’s performance. This discrepancy may stem from excessive averaging in EMA, where $\boldsymbol{\theta}^{*}_{t}$ is modified multiple times within the same task, potentially leading to suboptimal outcomes. Differently, our method applies weight averaging only after completing each task, enhancing computational efficiency by selectively focusing on pertinent model weights while retaining previous knowledge.
@@
Effect of PTMs. In this section, we evaluate the performance of the CoFiMA approach across a variety of backbone architectures, including self-supervised (MAE [17], MoCoV3 [5], and DINOv2 [45]) and supervised (ViT-Tiny [27] and ViT-B/16-SAM [13]) models. This comprehensive analysis aims to ascertain the adaptability and performance consistency of CoFiMA in diverse training paradigms. Results are visualized in Fig. 4.
@@
Our results indicate that CoFiMA enhances performance across almost all tested backbones relative to the baseline SLCA method. For instance, with the ViT-Tiny backbone, CoFiMA improves the Last-Accuracy of SLCA from 80.25% to 82.96%. This trend is similarly observed with ViT-Large, where CoFiMA achieves an accuracy of 86.81%, surpassing SLCA’s 85.93%. Only in the case of ViT-B/16-DINOv2, there is a slight reduction in performance, likely due to the benchmark reaching its saturation point: both SLCA and CoFiMA exhibit performance akin to the joint-learning gold standard.
@@
In the context of self-supervised learning models, CoFiMA demonstrates its efficacy by outperforming SLCA on ViT-B/16-MAE and ViT-B/16-MoCoV3 backbones. Importantly, the ViT-B/16-SAM backbone [13] achieves the best performance among the evaluated models. This can be attributed to the effective generalization features ingrained in the SAM backbone, a result of being trained with the SAM optimizer. This optimizer is known for its capability to enhance model generalizability, which is reflected in the superior performance metrics observed in our experiments, as also mentioned in Mehta et al. [40] work. We also notice that self-supervised pre-training usually results in larger performance gaps between continual learning baselines and joint training. Especially for ViT-B/16-MAE, as noted in Zhang et al. [85] work, because joint training using MAE requires smaller updates to learn all tasks compared to incremental learning using SLCA or CoFiMA.
@@
When compared to joint training, a standard upper bound in CL settings, CoFiMA demonstrates competitive performance. Although joint training always leads, as seen with ViT-Large (94.45%) and ViT-B/16-SAM (91.87%), CoFiMA remains close, particularly with ViT-B/16-SAM, where it achieves 90.48%. These results indicate that CoFiMA is a versatile approach, effective in enhancing performance with various backbones in both supervised and self-supervised learning contexts. However, the performance boost from our approach varies with the choice of backbone (size) and its pre-training paradigm.